OpenStudy (anonymous):

6 over x+4=4+4 over x-2 what is the two Xs. please help fast.

OpenStudy (mertsj):

\[\frac{6}{x+4}=4+\frac{4}{x-2}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

what is the two Xs

OpenStudy (mertsj):

Solve the equation to find x. Begin by multiplying each term by the common denominator.

OpenStudy (anonymous):

I just need the answer please

OpenStudy (mertsj):

1/2, -2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how did you get that answer

OpenStudy (mertsj):

99 Mertsj 0Solve the equation to find x. Begin by multiplying each term by the common denominator